x -1 0 1

f(x) 1/8 1/2 2

The table above shows some values for the function f.If f(x)=ka^x for some constants k and a,what is the value of a?
A)1/2 B) 1/4 C) 2 D) 4 E) 16


Answer by ewatrrr(24785)


Hi
x -1 0 1
f(x) 1/8 1/2 2
f(x) = ka^x |Using ordered pair (0,1/2)
1/2 = k a^0
1/2 = k
f(x) = (1/2)a^x |using ordered pair (1,2)
2 = (1/2)a^1
4 = a